An electron is moving in a circular orbit of radius 12 cm in a uniform 0.30 T magnetic field perpendicular to the velocity of the electron. What is the linear speed of the electron in units of m/s? (me=9.11x10-31 kg, qe=1.6x10-19 C)

Lütfen birini seçin:

a. 3.16x1011

b. 3.16x109

c. 3.16x1010

d. 6.32x109

e. 6.32x1010
Answer #1

DATE // Solution given, B=0.30 I me = gilly 10-31 kg ge = 1.6410-19 C r=12cm = 12x10^2 m for this system. f- fo Centripital h